Calculate the voltage of a battery connected to a parallel plate capacitor

with a plate area of 2.00x10-2 m2 and a plate separation of 2.0x10-3 m if

the charge stored is 4.00x10-12 C.


1
Expert's answer
2022-05-24T13:56:37-0400
"V=q/C\\
=\frac{q*d}{\epsilon_0A }"

"V=\frac{4.00*10^{-12}*2.0*10^{-3}}{8.85*10^{-12}*2.00*10^{-2} }=4.52*10^{-2}\:\rm V"